# Non-explosive pre-supernova feedback in the COLIBRE model of galaxy formation

### v1

@@ Abstract
pubyear: 2025pagerange: Non-explosive pre-supernova feedback in the COLIBRE model of galaxy formation20
@@ 1 Introduction
Although there is broad agreement on the necessity of feedback, the question of how best to model it has received intense attention over the past two decades. Various strategies have been adopted to address the notorious artificial inefficiency of thermal feedback due to limited numerical resolution: decoupling gas particles from hydrodynamics while imparting kinetic kicks (e.g., Springel & Hernquist, 2003), delaying radiative cooling (e.g., Stinson et al., 2006, 2013), or injecting large amounts of thermal energy to ensure long cooling times (e.g., Booth & Schaye, 2009; Dalla Vecchia & Schaye, 2012). Despite their differences, these approaches can all reproduce realistic galaxy populations, but only if carefully calibrated to match observed benchmarks (see, e.g., Crain & van de Voort, 2023, for a recent review).
@@
At low resolution, unresolved collapse and the imposition of an effective interstellar medium (ISM) equation of state (e.g. Springel & Hernquist, 2003; Schaye & Dalla Vecchia, 2008) force star formation to occur in a relatively smooth medium. In contrast, at high resolution or when the effective equation of state is dropped, stars emerge in dense, clustered, self-gravitating clouds. This transition alters galaxy properties profoundly. Star formation may proceed too rapidly within individual clouds due to short freefall times, while SN feedback can become ineffective if explosions occur in dense gas that cools rapidly, dissipating the injected energy. Moreover, if cloud collapse times are shorter than the lifetimes of massive stars, SNe may explode too late to regulate star formation occurring deep in the densest regions. These challenges make it difficult to achieve numerical convergence without careful (and expensive) recalibration.
@@ 2 The COLIBRE early feedback model
Figure 1: Top: Decomposition of the momentum rate, predicted by BPASS for a Chabrier (2003) IMF, into contributions from stellar winds (solid lines) and radiation pressure (dashed lines). While radiation pressure dominates the total momentum budget, only a fraction of this momentum couples effectively to the gas (see Fig. 2). Middle: Total specific momentum rate as a function of stellar age (x-axis) and metallicity (coloured lines, as indicated by the legend). Bottom: Cumulative total specific momentum as a function of age and metallicity.
@@ 2.1 Stellar winds and radiation pressure from young stars
To compute $\dot{p}_{\rm sw}$ for a given simple stellar population (SSP), we adopt the BPASS stellar population models (Eldridge et al., 2017; Stanway & Eldridge, 2018), specifically using the model suite that includes the effects of binary interactions, for the Chabrier (2003) stellar initial mass function with mass limits between $0.1$-$100\rm\ M_{\odot}$. BPASS provides the stellar wind-driven mass-loss rate, $\dot{m}_{\rm sw}$, and the associated energy output, $\dot{E}_{\rm sw}$, as a function of the SSP age and metallicity. Metallicities outside the tabulated range are clipped.
@@
For dust absorption, we adopt the Draine (2003) absorption cross section, $\sigma_{\rm dust}(\lambda)$, for the Weingartner & Draine (2001) carbonaceous-silicate grain model with $R_{V}=A_{V}/(A_{B}-A_{V})=3.1$, where $A_{V}$ is the total absorption in the $V$ band and $(A_{B}-A_{V})$ is the so-called selective absorption. The grain abundances have been renormalized by a factor of 0.93 to better match the local ISM, following Draine (2003). The dust mass per hydrogen nucleon is $m_{\rm dust/H}=1.398\times 10^{-26}\rm g\ H^{-1}$, yielding a dust-to-gas ratio $(D/G)_{\odot}=5.9\times 10^{-3}$ for the solar hydrogen mass fraction, $X_{\rm H}=0.74$. The corresponding dust optical depth, which depends on the local gas properties, is then:
@@
For the Lyman-Werner band (911.65 Å $\lambda<$ 1107 Å), which governs H2 self-shielding, we adopt the fitting function for the radiation suppression factor $S_{\mathrm{self}}^{\rm H_{2}}$ from CHIMES (equation 3.12 in Richings et al., 2014a). This depends on the gas temperature, $T$, the H2 column density, $N_{\rm H_{2}}$, and the turbulent velocity dispersion, $v_{\mathrm{turb}}$, which is assumed constant at $v_{\mathrm{turb}}=6\,\mathrm{km\,s}^{-1}$ following Ploeckinger et al. (2025). The corresponding optical depth is:
@@ 2.3.1 Momentum injection
In an SPH simulation, directly injecting momentum into dense gas particles is prone to rapid dissipation: because these particles are tightly packed and have short hydrodynamical timescales, the injected kinetic energy quickly thermalizes and is redistributed to neighboring particles through pressure forces and shocks, preventing the feedback from driving significant gas motions. Moreover, radiation pressure couples efficiently only to the densest gas, further limiting the impact of kinetic momentum injection. To circumvent these potential issues, we adopt a stochastic momentum injection scheme similar to that of Dalla Vecchia & Schaye (2008), in which the available momentum is delivered in discrete kicks with controlled velocities, ensuring that the total momentum budget is conserved while allowing individual particles to respond dynamically before dissipation dominates. These kicks are set to have a large velocity relative to the medium sound speed, which effectively causes the total available momentum to be distributed only to a subset of particles of the star SPH kernel. Note that here we refer to the SPH kernel of a stellar particle, rather than that of a gas particle. Although stellar particles are collisionless in COLIBRE, the code assigns them an effective SPH kernel that defines their domain of influence based on the surrounding gas particles. This kernel is used both for injecting stellar feedback and for continuously distributing metals into the surrounding gas. We leverage this infrastructure for our stochastic NEPS momentum injection implementation.
@@
We explore the impact of varying this parameter in Appendix A, where we demonstrate that the specific value of the kick velocity has only a minor impact on our results. Furthermore, as we show in Section 5.2, thermal feedback from H ii regions is the dominant regulatory process within our NEPS feedback model. The outsized importance of this thermal component renders the precise choice of the momentum kick velocity uncritical.
@@ 2.3.2 H ii regions
Our implementation is therefore functionally a “delayed cooling” scheme, which artificially sustains a high-pressure, photoionized phase around young stars for a duration comparable to the early lifetime of massive stars, until the onset of core-collapse SNe. This approach ensures that the thermal energy budget from photoionization is effectively coupled to the gas as mechanical work, providing the pressure support needed to regulate local star formation prior to the first core-collapse SN events begin at roughly 3.5 Myr. This method should be regarded as a subgrid model for the integrated pressure exerted by an H ii region over its early lifetime, rather than as a direct simulation of the ionization–recombination cycle around star-forming regions.
@@ 3.1.2 Star formation
$\frac{\sigma_{\rm therm}^{2}+\sigma_{\rm turb}^{2}}{GN_{\rm ngb}^{2/3}m_{\rm g,i}^{2/3}\rho_{\rm gas}^{1/3}}<1$ (21)
In the previous equation, $\sigma_{\rm therm}$ and $\sigma_{\rm turb}$ are the thermal and turbulent velocity dispersion, respectively; $m_{\rm g}$ and $\rho_{\rm gas}$ are, respectively, the mass of the $i$-th gas particle and its density, and $G$ is the gravitational constant.
@@ 3.1.3 Core-collapse supernova feedback
COLIBRE includes a SN feedback model designed to drive large-scale galactic winds. The scheme builds on the stochastic thermal feedback method of Dalla Vecchia & Schaye (2012), in which the energy from core-collapse SNe is accumulated and released in discrete heating events. In each event, neighbouring gas particles are heated by a large temperature increment, $\Delta T_{\rm SN}$, ensuring that the injected energy is not immediately radiated away.
@@ 4.1 Initial conditions
To test the implementation of our NEPS feedback module in COLIBRE, we consider redshift $z=0$ simulations of isolated gaseous disks embedded in an external Navarro-Frenk-White dark matter halo (Navarro et al., 1996) of virial [^2] We define virial quantities as those measured within a sphere where the mean density is 200 times the critical density of the Universe. mass, $M_{200}=1.5\times 10^{12}\ \rm M_{\odot}$, and concentration, $c=10$. To mimic the presence of an already formed thick stellar disk, we also include an external axisymmetric Miyamoto-Nagai potential (Miyamoto & Nagai, 1975, MN hereafter) with a mass, $M_{\rm MN}=3\times 10^{10}\rm\ M_{\odot}$, a scale radius, $R_{\rm MN}=4\rm\ kpc$, and a scale height, $Z_{\rm MN}=0.4\rm\ kpc$.
@@ 5.1 Simulations with gas cooling, heating, and star formation
Figure 10: Face-on surface density maps of the gas (left) and stellar (right) distribution at $t=500\rm\ Myr$. The simulations vary the gaseous disk mass, $M_{\rm d}$ (columns), and the gas particle mass resolution (rows). These runs include gas cooling and heating, star formation, and our NEPS feedback implementation. In all cases, supernova feedback is neglected. As in Fig. 7, the colour map assigns bluer colours to low-density pixels and redder colours to high-density pixels in a logarithmic scale, and the minimum and maximum density values shown are identical to those of Fig. 7.
@@ 5.2 Gas cooling and heating, star formation, and early feedback
Figure 11: Star formation rate as a function of time, measured in bins equally spaced every 10 $\rm Myr$. From top to bottom, the panels display the results for disks of different gas mass (as indicated), while the different colored lines show results for varying resolutions (as indicated by the legend in the middle panel). These simulations include gas cooling, star formation, and early feedback.
@@
Figure 12: As Fig. 9 but for simulations that include NEPS feedback. These distributions summarize the thermodynamic state of the gas in the absence of SN feedback but including our NEPS feedback. This figures highlights how NEPS feedback dramatically improves the convergence of the simulations.
@@
To quantify the mass contribution of the ionized phase, Fig. 14 shows the instantaneous mass fraction of gas in H ii regions, defined as particles with a temperature $T>10^{3.9}$ K. This figure confirms that although H ii regions are crucial for the regulation of star formation, they are not mass-dominant. They contribute less than 5105-10 % of the gas mass at any given density, and their fractional importance decreases at higher densities. [^5] These fractions are expected to decrease even further if the star formation rate decreases, which we expect to be the case if SN feedback is included.
@@
Figure 16: Face-on surface density maps of the gas distribution at $t=500\rm\ Myr$. The simulations vary the gaseous disk mass, $M_{\rm d}$ (columns), and the gas particle mass resolution (rows). The left panels (a) show simulations that include cooling, star formation, and supernova feedback. The right panels (b) show simulations that include cooling, star formation, our NEPS feedback module, and supernova feedback. A comparison reveals that NEPS feedback moderates the impact of SNe, leading to a more homogeneous ISM, particularly in high-resolution simulations. The color map and scale are identical to those in Fig. 7.
@@ 5.3 The inclusion of supernova feedback
In simulations lacking NEPS feedback, the disk is characterized by large, low-density cavities, seen mostly in blue colours. These structures are the result of highly clustered, energetic SN events occurring in the dense gas near the disk midplane. For instance, the high-resolution intermediate-mass gas disk shows several prominent “bubbles” that disrupt the outer disk. In contrast, when the NEPS module is active, the disk is visibly more homogeneous. SN-driven bubbles are systematically smaller and less numerous, indicating that the overall impact of SN feedback has been substantially mitigated.
@@
The differences across simulations with and without early feedback are a direct consequence of the regulatory effect of NEPS feedback on star formation. By increasing the pressure support within star-forming clouds, early feedback prevents the formation of overly dense stellar clusters where star formation would otherwise proceed with high efficiency. This leads to a less clustered, more spatially, and temporally distributed population of young stars. As a result, subsequent supernova explosions are less correlated, preventing the highly concentrated energy injection required to create large-scale bubbles and stir the disk.

### v2

@@ Abstract
pubyear: 2026pagerange: Non-explosive pre-supernova feedback in the COLIBRE model of galaxy formation21
@@ 1 Introduction
Although there is broad agreement on the necessity of feedback, the question of how best to model it has received intense attention over the past two decades. Various strategies have been adopted to address the notorious artificial inefficiency of thermal feedback due to limited numerical resolution: decoupling gas particles from hydrodynamics while imparting kinetic kicks (e.g., Springel and Hernquist, 2003), delaying radiative cooling (e.g., Stinson et al., 2006, 2013), or injecting large amounts of thermal energy to ensure long cooling times (e.g., Booth and Schaye, 2009; Dalla Vecchia and Schaye, 2012). Despite their differences, these approaches can all reproduce realistic galaxy populations, but only if carefully calibrated to match observed benchmarks (see, e.g., Crain and van de Voort, 2023, for a recent review).
@@
At low resolution, unresolved collapse and the imposition of an effective interstellar medium (ISM) equation of state (e.g. Springel and Hernquist, 2003; Schaye and Dalla Vecchia, 2008) force star formation to occur in a relatively smooth medium. In contrast, at high resolution or when the effective equation of state is dropped, stars emerge in dense, clustered, self-gravitating clouds. This transition alters galaxy properties profoundly. Star formation may proceed too rapidly within individual clouds due to short freefall times, while SN feedback can become ineffective if explosions occur in dense gas that cools rapidly, dissipating the injected energy. Moreover, if cloud collapse times are shorter than the lifetimes of massive stars, SNe may explode too late to regulate star formation occurring deep in the densest regions. These challenges make it difficult to achieve numerical convergence without careful (and expensive) recalibration.
@@ 2 The COLIBRE early feedback model
Figure 1: Top: Decomposition of the momentum rate, predicted by BPASS for a Chabrier (2003) IMF, into contributions from stellar winds (solid lines) and radiation pressure (dashed lines). While radiation pressure dominates the total momentum budget, only a fraction of this momentum couples effectively to the gas (see Fig. 2). Middle: Total specific momentum rate as a function of stellar age (x-axis) and metallicity (coloured lines, as indicated by the legend). Bottom: Cumulative total specific momentum as a function of age and metallicity. For comparison, the dashed line indicates the cumulative momentum rate expected from core-collapse SN in COLIBRE.
@@ 2.1 Stellar winds and radiation pressure from young stars
To compute $\dot{p}_{\rm sw}$ for a given simple stellar population (SSP), we adopt the BPASS stellar population models (Eldridge et al., 2017; Stanway and Eldridge, 2018), specifically using the model suite that includes the effects of binary interactions, for the Chabrier (2003) stellar initial mass function with mass limits between $0.1$-$100\rm\ M_{\odot}$. BPASS provides the stellar wind-driven mass-loss rate, $\dot{m}_{\rm sw}$, and the associated energy output, $\dot{E}_{\rm sw}$, as a function of the SSP age and metallicity. Metallicities outside the tabulated range are clipped.
@@
For dust absorption, we adopt the Draine (2003) absorption cross section, $\sigma_{\rm dust}(\lambda)$, for the Weingartner and Draine (2001) carbonaceous-silicate grain model with $R_{V}=A_{V}/(A_{B}-A_{V})=3.1$, where $A_{V}$ is the total absorption in the $V$ band and $(A_{B}-A_{V})$ is the so-called selective absorption. The grain abundances have been renormalized by a factor of 0.93 to better match the local ISM, following Draine (2003). The dust mass per hydrogen nucleon is $m_{\rm dust/H}=1.398\times 10^{-26}\rm g\ H^{-1}$, yielding a dust-to-gas ratio $(D/G)_{\odot}=5.9\times 10^{-3}$ for the solar hydrogen mass fraction, $X_{\rm H}=0.74$. The corresponding dust optical depth, which depends on the local gas properties, is then:
@@
For the Lyman-Werner band (911.65 Å $\lambda<$ 1107 Å), which governs H2 self-shielding, we adopt the fitting function for the radiation suppression factor $S_{\mathrm{self}}^{\rm H_{2}}$ from CHIMES (equation 3.12 in Richings et al., 2014a). This depends on the gas temperature, $T$, the H2 column density, $N_{\rm H_{2}}$, and the turbulent velocity dispersion, $\sigma_{\mathrm{turb,1D}}$, which is assumed constant at $\sigma_{\mathrm{turb,1D}}=6\,\mathrm{km\,s}^{-1}$ following Ploeckinger et al. (2025). The corresponding optical depth is:
@@ 2.3.1 Momentum injection
In an SPH simulation, directly injecting momentum into dense gas particles is prone to rapid dissipation: because these particles are tightly packed and have short hydrodynamical timescales, the injected kinetic energy quickly thermalizes and is redistributed to neighboring particles through pressure forces and shocks, preventing the feedback from driving significant gas motions. Moreover, radiation pressure couples efficiently only to the densest gas, further limiting the impact of kinetic momentum injection. To circumvent these potential issues, we adopt a stochastic momentum injection scheme similar to that of Dalla Vecchia and Schaye (2008), in which the available momentum is delivered in discrete kicks with controlled velocities, ensuring that the total momentum budget is conserved while allowing individual particles to respond dynamically before dissipation dominates. These kicks are set to have a large velocity relative to the medium sound speed, which effectively causes the total available momentum to be distributed only to a subset of particles of the star SPH kernel. Note that here we refer to the SPH kernel of a stellar particle, rather than that of a gas particle. Although stellar particles are collisionless in COLIBRE, the code assigns them an effective SPH kernel that defines their domain of influence based on the surrounding gas particles. This kernel is used both for injecting stellar feedback and for continuously distributing metals into the surrounding gas. We leverage this infrastructure for our stochastic NEPS momentum injection implementation.
@@
We explore the impact of varying this parameter in Appendix A, where we demonstrate that for momentum to be effective at suppressing runaway collapse and regulating the star formation history of an unstable disk, the specific value of the kick velocity needs to be highly supersonic, with velocities exceeding $\sim 20\rm\ km\ s^{-1}$. For stable disks, the exact choice of the kick velocity is largely irrelevant. Furthermore, as we show in Section 5.2, thermal feedback from H ii regions is the dominant regulatory process within our NEPS feedback model.
@@ 2.3.2 H ii regions
Our implementation is therefore functionally a “delayed cooling” scheme, which artificially sustains a high-pressure, photoionized phase around young stars for a duration comparable to the early lifetime of massive stars, until the onset of core-collapse SNe. This approach ensures that the thermal energy budget from photoionization is effectively coupled to the gas as mechanical work, providing the pressure support needed to regulate local star formation prior to the first core-collapse SN events begin at roughly 3 Myr (see the dashed line in the bottom panel of Fig. 1). This method should be regarded as a subgrid model for the integrated pressure exerted by an H ii region over its early lifetime, rather than as a direct simulation of the ionization–recombination cycle around star-forming regions.
@@ 3.1.2 Star formation
$\frac{\sigma_{\rm therm}^{2}+\sigma_{\rm turb,3D}^{2}}{GN_{\rm ngb}^{2/3}m_{\rm g,i}^{2/3}\rho_{\rm gas}^{1/3}}<1$ (21)
In the previous equation, $\sigma_{\rm therm}$ and $\sigma_{\rm turb,3D}$ are the thermal and 3D turbulent velocity dispersion, respectively; $m_{\rm g}$ and $\rho_{\rm gas}$ are, respectively, the mass of the $i$-th gas particle and its density, and $G$ is the gravitational constant.
@@ 3.1.3 Core-collapse supernova feedback
COLIBRE includes a SN feedback model designed to drive large-scale galactic winds. The scheme builds on the stochastic thermal feedback method of Dalla Vecchia and Schaye (2012), in which the energy from core-collapse SNe is accumulated and released in discrete heating events. In each event, neighbouring gas particles are heated by a large temperature increment, $\Delta T_{\rm SN}$, ensuring that the injected energy is not immediately radiated away.
@@ 4.1 Initial conditions
To test the implementation of our NEPS feedback module in COLIBRE, we consider redshift $z=0$ simulations of isolated gaseous disks embedded in an external Navarro-Frenk-White dark matter halo (Navarro et al., 1996) of virial [^2] We define virial quantities as those measured within a sphere where the mean density is 200 times the critical density of the Universe. mass, $M_{200}=1.5\times 10^{12}\ \rm M_{\odot}$, and concentration, $c=10$. To mimic the presence of an already formed thick stellar disk, we also include an external axisymmetric Miyamoto-Nagai potential (Miyamoto and Nagai, 1975, MN hereafter) with a mass, $M_{\rm MN}=3\times 10^{10}\rm\ M_{\odot}$, a scale radius, $R_{\rm MN}=4\rm\ kpc$, and a scale height, $Z_{\rm MN}=0.4\rm\ kpc$.
@@ 5.1 Simulations with gas cooling, heating, and star formation
Figure 10: Face-on surface density maps of the gas (left) and stellar (right) distribution at $t=500\rm\ Myr$. The simulations vary the gaseous disk mass, $M_{\rm d}$ (columns), and the gas particle mass resolution (rows). These runs include gas cooling and heating, star formation, and our NEPS feedback implementation. In all cases, supernova feedback is neglected. As in Fig. 7, the color map is idential to that of Fig. 7.
@@ 5.2 Gas cooling and heating, star formation, and early feedback
Figure 11: Star formation rate as a function of time, measured in bins equally spaced every 10 $\rm Myr$. From top to bottom, the panels display the results for disks of different gas mass (as indicated), while the different colored lines show results for varying resolutions (as indicated by the legend in the middle panel). These simulations include gas cooling, star formation, and early feedback. For comparison, the dashed lines in each panel show the feedback-free baseline at the highest resolution for each disk, corresponding to the results in Fig. 8.
@@
Figure 12: As Fig. 9 but for simulations that include NEPS feedback (solid lines). These distributions summarize the thermodynamic state of the gas in the absence of SN feedback but including our NEPS feedback. This figures highlights how NEPS feedback dramatically improves the convergence of the simulations. For comparison, the dashed lines show the results for the feedback-free simulations.
@@
To quantify the mass contribution of the ionized phase, Fig. 14 shows the instantaneous mass fraction of gas in H ii regions, defined as particles with a temperature $T>10^{3.9}$ K. This figure confirms that although H ii regions are crucial for the regulation of star formation, they are not mass-dominant. They contribute less than 5105-10 % of the gas mass at any given density, and their fractional importance decreases at higher densities. [^5] These fractions are expected to decrease even further if the star formation rate decreases, which we expect to be the case if SN feedback is included. While the precise mass fraction in H ii regions is influenced by our choice of $\Delta t_{\rm HII}$, the qualitative result—that star formation is regulated by a phase that occupies a small fraction of the total gas mass—is physically consistent with the small filling factor of compact H ii regions in dense molecular clouds (e.g. Kennicutt, 1984).
@@
Figure 16: Face-on surface density maps of the gas distribution at $t=500\rm\ Myr$. The simulations vary the gaseous disk mass, $M_{\rm d}$ (columns), and the gas particle mass resolution (rows). The left panels (a) show simulations that include cooling, star formation, and supernova feedback. The right panels (b) show simulations that include cooling, star formation, our NEPS feedback module, and supernova feedback. A comparison reveals that NEPS feedback moderates the impact of SNe, leading to a more homogeneous ISM, particularly in high-resolution simulations. The color map is identical to that of Fig. 7.
@@
We end by noting that while thermal feedback from H ii regions is the single most dominant mechanism to prevent initial runaway collapse of the disk, its efficacy can diminish at later times ($t>300$ Myr) in high-resolution simulations (e.g., the m4 run of the $10^{10}\,{\rm M}_{\odot}$ disk). In these cases, without momentum feedback, the gas can eventually settle into high-density configurations where recombination rates are high, reducing the duty cycle of thermal pressurization. However, the hierarchy between processes remains clear. Simulations lacking thermal feedback consistently undergo catastrophic fragmentation and starbursts, whereas those lacking momentum feedback remain largely regulated, albeit with higher variations in their star formation histories. Thus, we identify the thermal channel as the primary regulator that enables the system to reach a stable state, while the momentum channels help regulate that state and prevent the gas secular drift toward higher densities.
@@ 5.3 The inclusion of supernova feedback
In simulations lacking NEPS feedback, the disk is characterized by large, low-density cavities, seen mostly in blue colours. These structures are the result of highly clustered, energetic SN events occurring in the dense gas near the disk midplane. For instance, the high-resolution intermediate-mass gas disk shows several prominent “bubbles” that disrupt the outer disk. In contrast, when the NEPS module is active, the disk is visibly more homogeneous. SN-driven bubbles are systematically smaller and less numerous, indicating that the morphological disruption by SN feedback has been mitigated. However, the global regulation of star formation is enhanced, as shown by the lower SFR in Fig. 19.
@@
The differences across simulations with and without early feedback are a direct consequence of the regulatory effect of NEPS feedback on star formation. By increasing the pressure support within star-forming clouds, early feedback prevents the formation of overly dense stellar clusters where star formation would otherwise proceed with high efficiency. This leads to a less clustered, more spatially, and temporally distributed population of young stars. As a result, subsequent supernova explosions are less correlated, preventing the highly concentrated energy injection required to create large-scale bubbles and stir the disk. [^6] While this effect is qualitatively evident in the gas morphology and low-density bubbles distribution, a rigorous quantification would require, for instance, calculating the two-point correlation function of supernova events or young stars.